# PII-Bench: Evaluating Query-Aware Privacy Protection Systems

### v1

@@ Abstract
The widespread adoption of Large Language Models (LLMs) has raised significant privacy concerns regarding the exposure of personally identifiable information (PII) in user prompts. To address this challenge, we propose a query-unrelated PII masking strategy and introduce PII-Bench, the first comprehensive evaluation framework for assessing privacy protection systems. PII-Bench comprises 2,842 test samples across 55 fine-grained PII categories, featuring diverse scenarios from single-subject descriptions to complex multi-party interactions. Each sample is carefully crafted with a user query, context description, and standard answer indicating query-relevant PII. Our empirical evaluation reveals that while current models perform adequately in basic PII detection, they show significant limitations in determining PII query relevance. Even state-of-the-art LLMs struggle with this task, particularly in handling complex multi-subject scenarios, indicating substantial room for improvement in achieving intelligent PII masking.
{CJK}
UTF8gbsn
@@ 1 Introduction
Current practices reveal that the vast majority of users adopt a zero-protection approach when utilizing LLM services, submitting original prompts containing PII directly to the LLMs. While an obvious protection strategy would be to mask all PII ( Nakamura et al., 2020Biesner et al., 2022;Lukas et al., 2023), as shown in Fig. 1, this approach significantly compromises service quality. An ideal Privacy Protection System should maintain LLMs’ functionality while maximizing user privacy protection. For instance, when a user inquires about a candidate’s suitability for a senior researcher position, masking their educational background and work experience would render the LLM incapable of making an effective assessment.
@@
The implementation of query-unrelated PII masking stragety faces two-tier challenges. The first involves accurate identification of all PII within the prompt, serving as foundational work. The second requires determining the relevance of identified PII to user queries. While existing research has made progress in basic PII detection, systematic studies considering query relevance remain scarce.
To advance the field of privacy-preserving language models, we present PII-Bench, a comprehensive evaluation framework designed to assess Privacy Protection Systems’ efficacy in preserving Large Language Models’ core functionalities while optimizing user privacy safeguards. PII-Bench comprising 2,842 carefully designed test samples across 55 fine-grained PII categories, ranging from basic personal information to complex social relationship data. Each sample consists of three key components: (1) A user query simulating real-world information needs. (2) A context description containing diverse PII. (3) A standard answer indicating query-relevant PII and masking requirements.
Our experimental analysis reveals that while existing models, including Bidirectional Long Short-Term Memory with Conditional Random Fields (BiLSTM-CRF), perform adequately in basic PII detection, they demonstrate notable limitations in determining PII query relevance. Even state-of-the-art LLMs face challenges in this task, indicating substantial room for improvement in achieving intelligent PII masking. Despite the recent advances in model architecture and training techniques, smaller models (SLM) still show considerable performance gaps compared to larger LLMs, particularly in determining PII query relevance.
@@ 3.2 PII-Bench Construction
Based on the task definition above, we designed an automated process for constructing the PII evaluation dataset, as illustrated in Fig. 2.
@@ 3.2.1 PII Entity Generation
Following Papadopoulou et al. (2022), we expanded the PII type set $\mathcal{T}$ into 55 subcategories (see Appendix A), employing two complementary strategies for entity generation:
(1) Rule-based Generation: Applicable for deterministic PII types with fixed formats or enumerable value sets, such as phone numbers, email addresses, and standardized ID numbers. (2) LLM-based Generation: Applicable for non-deterministic PII types requiring contextual understanding and real-world knowledge, such as occupation descriptions and detailed addresses. This method leverages GPT-4-0806 to generate semantically appropriate and contextually relevant entities.
@@ 3.2.2 User Description Generation
(1) Entity Selection: For subject $s$, randomly sample $n$ entities ($4\leq n\leq 16$) from different PII types to construct entity set $\mathcal{E}$. The sampling process ensures diversity of PII types while considering their natural distribution in real-world scenarios. (2) Consistency Optimization: Ensure logical compatibility among entities in $\mathcal{E}$ through designed verification rules. For example, verifies reasonable correspondence between age and educational history as shown in Fig. 2. (3) User Desc Generation: Selects appropriate expression styles to generate the user description. It employs formal description formats like job resumes and employee records in professional scenarios; casual expressions like personal profiles and self-introductions in social scenarios.
@@
(1) Entity Selection: Construct relationship network $R(s_{i},s_{j})$ for subject pairs $(s_{i},s_{j})$. Relationship types include intersection relationships like colleagues and alumni, hierarchical relationships like parent-child and teacher-student, and non-intersection relationships with no direct connection. (2) Consistency Optimization: This stage first establishes entity dependency rules based on relationship type $R$. Then ensures consistency of shared attributes among related subjects, such as company address for employees of the same company. This stage also derives related attributes based on relationship type, such as age differences in parent-child relationships. And finally remove the sample which contains contradictions. (3) User Desc Generation: This stage designs natural interaction environments matching relationship characteristics, placing subjects in realistic scenarios (like meetings, family activities) and constructing multi-party dialogue flows to reflect interactive relationships.
@@ 3.2.3 Query Construction
For each description $d$, query construction follows a four-phase process:
(1) Entity Selection: Randomly sample $k$ entities ($1\leq k\leq 3$) from $\mathcal{E}$ to form query-relevant entity set $\mathcal{E}_{q}$. (2) Scenario Design: Construct query contexts that align with real-world application scenarios. The goal is to simulate actual user needs for PII information in specific situations. For example, when $\mathcal{E}_{q}$ contains “Work Experience”: “5 years as Machine Learning Engineer”, “Education Background”: “Stanford University Ph.D. in Computer Science”, this stage generates query scenarios like “As a hiring manager, I need to verify if this candidate’s education and relevant work experience meet the requirements for the Senior Researcher position”. (3) Entity Abstraction: Map specific PII entities in $\mathcal{E}_{q}$ to abstract representations, maintaining basic semantic properties while enhancing privacy protection. (4) Query Generation: Integrate abstract entities into corresponding scenarios through GPT-4-0806 model to generate natural queries $q$ that fit practical application scenarios.
@@ 3.2.4 Human Verification
All content generated by GPT-4-0806 undergoes rigorous verification by five professional annotators and the authors, focusing on: (1) Completeness and accuracy of PII entity annotations in description $d$. (2) Correspondence between query $q$ and query-relevant entity set $\mathcal{E}_{q}$. (3) Overall semantic coherence and scenario authenticity. Complete annotation guidelines and quality control procedures are detailed in Appendix E.
@@ 3.3 Dataset Partitioning and Statistics
Table 3 presents the partition and key statistics of PII-Bench, which comprises two main datasets (PII-single and PII-multi) and two specialized test sets (PII-hard and PII-distract). Each sample follows a consistent JSON structure containing four key components: user description, query, comprehensive PII entity annotations, and query-relevant PII labels, as illustrated in Fig. 3.
PII-Single and PII-Multi: Based on the number of subjects in descriptions, the dataset is divided into two main subsets. PII-Single contains 2000 description-query pairs involving single subjects, focusing on model performance in handling individual information. PII-Multi contains 2000 description-query pairs involving multiple related subjects, evaluating model capability in handling privacy information within complex interpersonal networks.
@@
Test-Distract Construction: Construct 200 samples simulating complex multi-user interaction scenarios. Each sample integrates five different descriptions $\{d_{1},...,d_{5}\}$ from PII-Single and PII-Multi, and constructs queries $q$ involving three of these descriptions based on professional networks, knowledge platforms, and community forum interaction templates. The generation process employs specific dialogue flow transformation strategies to ensure natural transitions and semantic coherence between multiple descriptions. Scenario design particularly emphasizes simulating real-world information interference and complex interaction patterns.
@@ 3.4 Human Performance
To establish a human baseline for PII-Bench, we recruited 25 graduate students specializing in data security from top universities across China. All participants had at least two years of research experience in privacy protection and information security. Before the formal evaluation, participants completed a comprehensive training session and passed a qualification test (detailed in Appendix C).
We designed two evaluation sets: a main test set comprising 400 randomly sampled instances (200 each from PII-single and PII-multi), and a challenging set of 100 instances from PII-distract. Each instance underwent independent assessment by five participants through our online evaluation platform. Participants performed two sequential tasks: PII recognition, which involved determining minimal text spans, associated subjects, and PII types for all entities in the user description, followed by query-relevant PII detection to identify entities essential for addressing the given query. The result of the human baseline is shown in Table 2.
@@
Table 2: Human performance in PII-Bench. Desc-F1 measures accuracy in the PII recognition task while Query-F1 evaluates the query-relevant PII detection task.
@@
Figure 3: An example from PII-Bench, which aims to evaluate Privacy Protection System’s ability by masking maximize PII while maintain LLM’s functionality. The evaluation is seperated by two fundamental tasks: (a) The PII Detection Task: Identify and classify PII entities for each subject in the prompt, with ground truth labels shown on the right side. (b) The Query-Related PII Detection Task: Determine which PII entities are necessary for answering the user query, enabling selective masking of irrelevant personal information.
@@ 4.1 Overall Setup
SLM Baselines: To investigate scaling effects, we included two small-scale language models: Llama-3.1-8B-Instruct (Llama3.1-SLM) and Qwen-2.5-7B-Instruct (Qwen2.5-SLM). All experiments utilized default parameters with temperature set to 0 to ensure reproducibility.
Prompt Baselines: The assessment incorporated multiple prompting strategies for query-related PII detection. Naive inputs the user description and query. Naive /w Choice includes a list of candidate PII entities to constrain the selection space. Self-CoT (Wei et al., 2022) incorporating step-by-step reasoning prompts. Auto-CoT Zhang et al. (2022), which automates the generation of chain-of-thought demonstrations through three-shot setting. Self-Consistency (SC) Wang et al. (2022), which synthesizes multiple reasoning paths to derive the final output. Plan-and-Solve CoT (PS-CoTWang et al. (2023) develops a strategic plan before executing the solution process. Appendix D.3 provides details of each prompts.
Metrics: The PII detection task evaluates model performance through two sets of metrics: Strict-F1 measures the accuracy of subject identification, entity span detection, and PII type classification simultaneously. Ent-F1 focuses on entity span detection independent of subject attribution and type classification. For query-related detection, model performance is measured through Precision, Recall, and F1. Considering the inherent variation in entity expressions and potential partial matches, RougeL-F is employed for both tasks to complement the exact matching metrics. Detailed computation procedures are provided in Appendix D.1.
@@
Figure 5: Performance comparison across different models for seven main PII types.
@@ 4.2 Performance on Query-Unrelated PII Masking
We evaluate models’ performance on the query-unrelated PII masking task, which requires both accurate PII detection and relevance assessment. Table 4 presents our experimental results:
Joint task yields improved performance. Notably, models achieve higher F1 scores in this combined task compared to individual query-relevance tasks. GPT4o reaches 0.77 F1 with Self-Consistency prompting, suggesting that the joint objective may provide complementary signals. Open-source models demonstrate comparable capabilities, with both Llama3.1 and Qwen2.5 achieving 0.76 F1 using Auto-CoT.
@@ 4.3 Performance on PII Detection
Experimental results on the PII detection task, presented in Table 6, reveal several key findings:
Large Language Models demonstrate superior detection capabilities. API-based LLMs achieve strong performance across standard datasets, with DeepSeekV3 and GPT4o leading in Strict-F1 scores (0.903 and 0.891 on PII-Single and PII-Multi, respectively). The open-source Llama3.1 shows competitive performance, particularly in entity recognition (Ent-F1: 0.942 on PII-Multi), while traditional BiLSTM-CRF maintains reasonable entity detection capabilities despite its simpler architecture.
Entity type classification remains challenging. A consistent gap between Strict-F1 and Ent-F1 scores indicates that accurate PII type classification poses greater challenges than entity boundary detection. This disparity becomes more pronounced in the PII-Distract dataset, where models maintain relatively high Ent-F1 scores despite significant drops in Strict-F1, suggesting increased difficulty in precise PII categorization under complex scenarios.
@@ 4.4 Performance on Query-Related PII Detection
Tables 7 presents the results on PII-single dataset across different model scales and prompting strategies.
Limited Performance of Current LLMs. State-of-the-art LLMs exhibit limited performance in this task, with GPT4o achieving only 0.627 F1 score with naive prompting—substantially below human performance (0.951 F1). Open-source alternatives demonstrate competitive performance, with Qwen2.5 reaching 0.615 F1.
Impact of Advanced Prompting. Chain-of-thought approaches generally improve performance, with Self-Consistency and Auto-CoT proving most effective for different models (0.716 F1 for GPT4o with Self-Consistency; 0.710 F1 for Qwen2.5 with Auto-CoT). However, these benefits are highly dependent on model scale—smaller models often show degraded performance with complex prompting strategies, indicating insufficient reasoning capabilities.
Effectiveness of Entity Candidates. The provision of candidate PII entities (Naive w/ Choice) substantially improves performance across all models (e.g., GPT4o improves from 0.627 to 0.842 F1). However, the practical applicability of this approach is limited, as candidate entities are rarely available in real-world scenarios.

### v2

@@ Abstract
The widespread adoption of Large Language Models (LLMs) has raised significant privacy concerns regarding the exposure of personally identifiable information (PII) in user prompts. To address this challenge, we propose a query-unrelated PII masking strategy and introduce PII-Bench, the first comprehensive evaluation framework for assessing privacy protection systems. PII-Bench comprises 2,842 test samples across 7 PII types with 55 fine-grained subcategories, featuring diverse scenarios from single-subject descriptions to complex multi-party interactions. Each sample is carefully crafted with a user query, context description, and standard answer indicating query-relevant PII. Our empirical evaluation reveals that while current models perform adequately in basic PII detection, they show significant limitations in determining PII query relevance. Even advanced LLMs struggle with this task, particularly in handling complex multi-subject scenarios, indicating substantial room for improvement in achieving intelligent PII masking.
@@ 1 Introduction
Current practices reveal that the vast majority of users adopt a zero-protection approach when utilizing LLM services, submitting original prompts containing PII directly to the LLMs. While an obvious protection strategy would be to mask all PII ( Nakamura et al., 2020Biesner et al., 2022;Lukas et al., 2023), as shown in Figure 1, this approach significantly compromises service quality. An ideal Privacy Protection System should maintain LLMs’ functionality while maximizing user privacy protection. For instance, when a user inquires about a candidate’s suitability for a senior researcher position, masking their educational background and work experience would render the LLM incapable of making an effective assessment.
@@
The implementation of query-unrelated PII masking strategy faces two-tier challenges. The first involves accurate identification of all PII within the prompt, serving as foundational work. The second requires determining the relevance of identified PII to user queries. While existing research has made progress in basic PII detection, systematic studies considering query relevance remain scarce.
To advance the field of privacy-preserving language models, we present PII-Bench, a comprehensive evaluation framework designed to assess Privacy Protection Systems’ efficacy in preserving LLMs’ core functionalities while optimizing user privacy safeguards. PII-Bench comprising 2,842 carefully designed test samples across 7 PII types with 55 fine-grained subcategories, ranging from basic personal information to complex social relationship data. Each sample consists of three key components: (1) A user query simulating real-world information needs. (2) A context description containing diverse PII. (3) A standard answer indicating query-relevant PII and masking requirements.
Our experimental analysis reveals that while existing models, including Bidirectional Long Short-Term Memory with Conditional Random Fields (BiLSTM-CRF) Chen et al. (2017), perform adequately in basic PII detection, they demonstrate notable limitations in determining PII query relevance. Even advanced LLMs face challenges in this task, indicating substantial room for improvement in achieving intelligent PII masking. Despite the recent advances in model architecture and training techniques, Small Language Models (SLM) still show considerable performance gaps compared to larger LLMs, particularly in determining PII query relevance.
@@ 3.2 PII-Bench Construction
Based on the task definition above, we designed an automated process for constructing the PII evaluation dataset, as illustrated in Figure 2.
@@ 3.2.1 PII Entity Generation
Following Papadopoulou et al. (2022), we expanded the PII type set $\mathcal{T}$ with 7 main types into 55 fine-grained subcategories as detailed in Table 10, employing two complementary strategies for entity generation:
(1) Rule-based Generation: Applicable for deterministic PII types with fixed formats or enumerable value sets (e.g., phone numbers, email addresses, ID numbers). (2) LLM-based Generation: Applicable for non-deterministic PII types requiring contextual understanding and real-world knowledge (e.g., occupation descriptions, detailed addresses).
@@ 3.2.2 User Description Generation
(1) Entity Selection: For subject $s$, randomly sample $n$ entities ($4\leq n\leq 16$) from different PII types to construct entity set $\mathcal{E}$. The sampling process ensures diversity of PII types while considering their natural distribution in real-world scenarios. (2) Consistency Optimization: Ensure logical compatibility among entities in $\mathcal{E}$ through LLM-based verification with crafted prompts (detailed in Appendix C). For example, verifies reasonable correspondence between age and educational history as shown in Figure 2. (3) User Desc Generation: Selects appropriate expression styles to generate the user description. It employs formal description formats like job resumes and employee records in professional scenarios; casual expressions like personal profiles and self-introductions in social scenarios.
@@
(1) Entity Selection: Construct relationship network $R(s_{i},s_{j})$ for subject pairs $(s_{i},s_{j})$. Relationship types include intersection relationships like colleagues and alumni, hierarchical relationships like parent-child and teacher-student, and non-intersection relationships with no direct connection. (2) Consistency Optimization: Apply relationship-aware verification based on $R$ to maintain cross-subject coherence (detailed in Appendix C). The optimization enforces shared attributes for intersection relationships (e.g., company address for colleagues), structural constraints for hierarchical relationships (e.g., age differences for parent-child pairs), and removes samples with irresolvable contradictions. (3) User Desc Generation: This stage designs natural interaction environments matching relationship characteristics, placing subjects in realistic scenarios (like meetings, family activities) and constructing multi-party dialogue flows to reflect interactive relationships.
@@ 3.2.3 Query Construction
For each description $d$, we construct queries through four phases:
(1) Entity Selection: Randomly sample $k$ entities ($1\leq k\leq 3$) from $\mathcal{E}$ to form query-relevant entity set $\mathcal{E}_{q}$. (2) Scenario Design: Generate domain-specific contexts aligned with real-world applications (detailed in Appendix D). For instance, given $\mathcal{E}_{q}$ containing work experience and educational background, we design scenarios such as recruitment evaluation or career planning. (3) Entity Abstraction: Transform specific entities in $\mathcal{E}_{q}$ into abstract representations while preserving semantic properties. (4) Query Generation: Synthesize natural queries $q$ by integrating abstracted entities into their corresponding scenarios.
@@ 3.2.4 Human Verification
All content generated by GPT-4-0806 undergoes rigorous verification by five professional annotators and the authors, focusing on: (1) Completeness and accuracy of PII entity annotations in description $d$. (2) Correspondence between query $q$ and query-relevant entity set $\mathcal{E}_{q}$. (3) Overall semantic coherence and scenario authenticity. Complete annotation guidelines and quality control procedures are detailed in Appendix G.
@@ 3.3 Dataset Partitioning and Statistics
Table 3 presents the partition and key statistics of PII-Bench, which comprises two main datasets (PII-single and PII-multi) and two specialized test sets (PII-hard and PII-distract). Each sample follows a consistent JSON structure containing four key components: user description, query, comprehensive PII entity annotations, and query-relevant PII labels, as illustrated in Figure 3.
PII-Single and PII-Multi: Based on the number of subjects in descriptions, the dataset is divided into two main subsets. PII-Single contains 1,214 description-query pairs involving single subjects, focusing on model performance in handling individual information. PII-Multi contains 1,228 description-query pairs involving multiple related subjects, evaluating model capability in handling privacy information within complex interpersonal networks.
@@
Test-Distract Construction: Construct 200 samples simulating complex multi-user interaction scenarios. Each sample integrates five different descriptions from PII-Single and PII-Multi, and constructs queries involving three of these descriptions based on professional networks, knowledge platforms, and community forum interaction templates. The generation process employs dialogue flow transformation strategies to ensure natural transitions and semantic coherence, simulating real-world information interference and complex interaction patterns.
@@ 3.4 Human Performance
To establish a human baseline, we recruited 25 graduate students with at least two years of research experience in privacy protection and data security from top Chinese universities. All participants completed comprehensive training and passed a qualification test before formal evaluation (details in Appendix E). We evaluated 400 randomly sampled instances (200 each from PII-single and PII-multi) and 100 instances from PII-distract, with each instance independently assessed by five participants. Participants performed two sequential tasks: PII detection, which involved determining minimal text spans, associated subjects, and PII types for all entities in the user description, followed by query-relevant PII detection to identify entities essential for addressing the given query. The result of the human baseline is shown in Table 2.
@@
Table 2: Human performance in PII-Bench. PII-F1 measures accuracy in the PII detection task while Query-F1 evaluates the query-relevant PII detection task.
@@
Figure 3: An example from PII-Bench, which aims to evaluate Privacy Protection System’s ability by masking maximize PII while maintaining LLM’s functionality. The evaluation is seperated by two fundamental tasks: (a) The PII Detection Task: Identify and classify PII entities for each subject in the prompt, with ground truth labels shown on the right side. (b) The Query-Related PII Detection Task: Determine which PII entities are necessary for answering the user query, enabling selective masking of irrelevant personal information.
@@ 4.1 Overall Setup
SLM Baselines: To investigate scaling effects, we included two small-scale language models: Llama-3.1-8B-Instruct (Llama3.1-SLM) and Qwen-2.5-7B-Instruct (Qwen2.5-SLM). All experiments utilized default parameters with temperature set to 0 to ensure reproducibility. Additionally, we conducted a comprehensive evaluation on smaller, deployment-ready models (0.5B-3B parameters) to assess their viability for on-device PII protection, with detailed results presented in Appendix F.3.
Prompt Baselines: The assessment incorporated multiple prompting strategies for query-related PII detection. Naive inputs the user description and query. Naive /w Choice includes a list of candidate PII entities to constrain the selection space. Self-CoT (Wei et al., 2022) incorporating step-by-step reasoning prompts. Auto-CoT Zhang et al. (2022), which automates the generation of chain-of-thought demonstrations through three-shot setting. Self-Consistency (SC) Wang et al. (2022), which synthesizes multiple reasoning paths to derive the final output. Plan-and-Solve CoT (PS-CoTWang et al. (2023) develops a strategic plan before executing the solution process. Appendix F.6 provides details of each prompts.
Metrics: The PII detection task evaluates model performance through two sets of metrics: Strict-F1 measures the accuracy of subject identification, entity span detection, and PII type classification simultaneously. Ent-F1 focuses on entity span detection independent of subject attribution and type classification. For query-related detection, model performance is measured through Precision, Recall, and F1. Considering the inherent variation in entity expressions and potential partial matches, RougeL-F is employed for both tasks to complement the exact matching metrics. Detailed computation procedures are provided in Appendix F.1. All reported model performance metrics are mean scores over test sets (temperature=0, top_k=1).
@@ 4.2 Performance on Query-Unrelated PII Masking
We evaluate models’ performance on the query-unrelated PII masking task, which requires both accurate PII detection and relevance assessment. Table 4 presents our results:
Joint task yields improved performance. Models achieve higher F1 scores in this combined task compared to individual query-relevance tasks. GPT4o reaches 0.77 F1 with Self-Consistency prompting, suggesting complementary signals from the joint objective. Open-source models demonstrate comparable capabilities, with both Llama3.1 and Qwen2.5 achieving 0.76 F1 using Auto-CoT.
@@ 4.3 Performance on PII Detection
Results on the PII detection task (Table 6) reveal:
Large Language Models demonstrate superior detection capabilities. API-based LLMs achieve strong performance, with DeepSeekV3 and GPT4o leading in Strict-F1 scores (0.903 and 0.891 on PII-Single and PII-Multi, respectively). Open-source Llama3.1 shows competitive performance, particularly in entity recognition (Ent-F1: 0.942 on PII-Multi).
Entity type classification remains challenging. A consistent gap between Strict-F1 and Ent-F1 scores indicates that accurate PII type classification poses greater challenges than entity boundary detection. This disparity becomes more pronounced in the PII-Distract dataset, suggesting increased difficulty in precise PII categorization under complex scenarios.
@@ 4.4 Performance on Query-Related PII Detection
Table 7 presents the results on PII-single dataset across different model scales and prompting strategies.
Limited Performance of Current LLMs. Advanced LLMs exhibit limited performance, with GPT4o achieving only 0.627 F1 score with naive prompting—substantially below human performance (0.951 F1). Open-source alternatives show competitive performance, with Qwen2.5 reaching 0.615 F1.
Impact of Advanced Prompting. Chain-of-thought approaches generally improve performance, with Self-Consistency and Auto-CoT proving most effective (0.716 F1 for GPT4o with Self-Consistency; 0.710 F1 for Qwen2.5 with Auto-CoT). However, these benefits are highly dependent on model scale—smaller models often show degraded performance with complex prompting strategies.
Effectiveness of Entity Candidates. Providing candidate PII entities (Naive w/ Choice) substantially improves performance across all models (e.g., GPT4o improves from 0.627 to 0.842 F1). However, practical applicability is limited as candidate entities are rarely available in real-world scenarios.
4.5 Privacy-Utility Tradeoff Analysis
We evaluated the tradeoff between privacy protection and query utility across different PII masking strategies
Experimental Setup. We selected 200 unique user descriptions from PII-Bench and applied three masking methods: (1) No Mask: Original text with all PII preserved; (2) All PII Mask: All detected PII entities replaced with their corresponding tags; (3) Query-unrelated PII Mask: Only query-irrelevant PII entities masked. For each method, we generated responses using GPT4o.
Metrics. The privacy-utility tradeoff is evaluated through two metrics: Privacy Score ($P$): Quantifies the proportion of PII successfully protected in the processed text. Utility Score ($U$): Combines semantic similarity and LLM evaluation for quality assessment between mask and unmasked responses. Detailed computation procedures are provided in Appendix F.2.
Results. Table 8 shows that Query-unrelated PII Mask achieves optimal balance between privacy protection ($P=0.83$) and utility preservation ($U=0.89$), outperforming other strategies in balanced score ($B=0.86$).
Method P U B
No Mask 0.00 1.00 0.50
All PII Mask 1.00 0.52 0.76
Query-unrelated PII Mask 0.83 0.89 0.86
Table 8: Privacy-utility tradeoff across different PII masking strategies. Balanced scores ($B$) combine both metrics with equal weights.
4.6 Real-World Dataset Validation
To validate the reliability of synthetic data, we constructed PII-Real, a dataset comprising 100 instances derived from publicly available profiles of 20 AI researchers with manually annotated PII entities and human-written queries. Evaluation on PII-Real demonstrates consistent alignment with synthetic data. Comprehensive experimental results and validation analysis are provided in Appendix F.4.